Suppose you want to purchase a $305,000 home, and you have the required $63,000 down payment in savings.

Complete the table below for a mortgage option presented by a mortgage broker.

Principal: $242,000

Interest Rate: 3.255%

Terms (years): 20

(Find) Payment:

(Find) Total of Payments over Life of Loan:

(Find) Total Interest Paid over Life of Loan:
Answer #1

The payment amount is calculated using the PMT as follows:

=PMT(rate,nper,pv)

=PMT(3.255%,20,-242000)

=16,652.10

Total of payments =16,652.10*20 = 333,042.06

Total interest=333,042.06-242000=91,042.06